A 10kg stone is suspended with a rope of breaking strength 30kg-wt. The minimum time in which the stone can be raised through a height of 10m is
1
Expert's answer
2019-10-04T12:52:46-0400

Mass of stone = "10\ kg"

Breaking strength of rope = "30\ kg\ wt=30g\ =300\ N"

Let the maximum allowed acceleration be "a"

So

Breaking strength "=300=mg+ ma=100+10a"

"10a=200\\So,\\a=20\ m\ s^{-2}"


"h=10\ m\\let\ \ t\ \ be\ time \ taken\\"

"u=0\\using\ equation\ of \ motion\\h=ut+\frac{at^2}{2}=10"

So,

"t=\sqrt\frac{20}{a}=\sqrt\frac{20}{20}= 1\ sec ond"